# Synthesizing Sun-as-a-star flare spectra from high-resolution solar observations

### v1

@@ Abstract
Aims. We produce and analyze synthetic full-disk spectra derived from 19 small area field-of-view optical observations of solar flares acquired by the Swedish 1-m Solar Telescope (SST) between 2011 and 2024. These are used to investigate what can and cannot be inferred about physical processes on the Sun from Sun-as-a-star observations.
Methods. The recently released Numerical Empirical Sun-as-a-Star Integrator (NESSI) code provides synthetic full-disk integrated spectral line emission based on smaller field-of-view input, accounting for center-to-limb variations and differential rotation. We use this code to generate pseudo-Sun-as-a-star spectra from the SST observations.
Results. We show that limited-area solar observations can be extrapolated to represent the entire disk accurately in a manner close to what is achievable with Sun-as-a-star telescopes. Additionally, we identify nine spectral features, four of which are caused by instrumental and data reduction effects. Primarily, we find a relation between the heliocentric angle of flares and the width of the excess emission left by them, a source of false positive CME-like signatures, as well as define an energy scaling law based on chromospheric line intensities which shows that the peak flare contrast roughly scales with the square root of the bolometric energy.
Conclusions. The presented method of creating pseudo-Sun-as-a-star observations from limited field-of-view solar observations allows for the accurate comparison of solar flare spectra with their stellar counterparts, while allowing for the detection of signals at otherwise unachievable noise levels.
@@ 1 Introduction
Solar flares and the corresponding coronal mass ejections (CMEs) are the primary drivers of the most severe disruptions to the heliosphere (Cliver et al., 2022). When aimed at the Earth, they can cause widespread damage to electronics, blackouts, and disruptions in communications (e.g., Cole, 2003; Miteva et al., 2023). The energies of these flares are typically between and erg (Fletcher et al., 2011), which is still two orders of magnitude lower than the most powerful stellar flares. These extreme events are believed to reach energies as high as or erg, depending on whether one adopts the assumption of blackbody radiation (Maehara et al., 2012) or hydrogen recombination continuum (Simões et al., 2024) for energy fitting. While solar flares leave only faint (Woods et al., 2006; Kretzschmar, 2011), or at times no detectable (e.g., Pietrow et al., 2024a), signatures in the total solar irradiance, they can dramatically outshine their host stars when occurring on intrinsically dimmer M dwarfs (e.g., Zhan et al., 2019). They are also visible as dimmings in the corona and transition region (e.g. Hou et al., 2025; Veronig et al., 2025).
@@
Although such extreme events have been observed on other stars for years (e.g., Pietras et al., 2022; Namizaki et al., 2023; Bruno et al., 2024), stellar flares have only recently attracted significant attention in the context of exoplanet research. In part as sources of noise and transit time variation (e.g., Davenport et al., 2019; Howard et al., 2023), but also as agents of change in planetary environments where it is believed that flares and their ejecta can permanently alter planetary atmospheres and thus their habitability (e.g., Pulkkinen, 2007; Airapetian et al., 2020; Temmer, 2021; Ilin et al., 2024).
@@
For this reason, it is important to understand stellar flares and what sets them apart from their solar counterparts. However, this is not straightforward as solar flares are typically studied in high spectral and spatial resolution, while the opposite is true for stellar flares which are primarily observed in white light, and only more recently with high-resolution spectrographs (e.g., Namekata et al., 2021; Martínez González et al., 2022; Namizaki et al., 2023). As solar flares imprint only faint (Woods et al., 2006; Kretzschmar, 2011) or non-detectable (Pietrow et al., 2024a) white-light fluctuations on a disk-integrated scale, direct comparisons have so far only been possible within numerical models (e.g., Allred et al., 2005; Druett et al., 2024). However, this is no longer the case when looking at spectral lines where even relatively weak flares for stellar standards can leave an imprint on a disk-integrated spectrum (see Fig. 1).
@@
Namekata et al. (2021, 2022), Otsu et al. (2022), and Xu et al. (2022) pioneered a pseudo-Sun-as-a-star analysis technique for flares and CMEs by studying line-profile evolutions of spatially integrated spectra of a small region around the event. Here, a difference was seen between the imprint of the flare on the spectra and its position on the disk. Such a relation has already been known to exist between UV and X-ray radiation for solar flares (Woods et al., 2006), and has recently been suggested to also exist for spectral lines in the visible by Yu et al. (2025) through 1D flare simulations using the RADYN code (Allred et al., 2005). If confirmed in observations, such an effect could prove to be a powerful technique for determining the disk position of these events.
@@
Follow-up studies with such Sun-as-a-star telescopes are limited by their signal-to-noise ratio (S/N) and are primarily sensitive to very strong flares. On the other hand, there is a large collection of flare observations taken with high spatial and spectral resolution in the solar community that have excellent S/N but a limited FOV. Such observations can be processed in a way similar to that of other pseudo-Sun-as-a-star studies to obtain more statistics on flare behavior in a disk-integrated setting. However, we propose to add an additional step to more closely approximate true Sun-as-a-star flare observations. In this work, we apply the Numerical Empirical Sun-as-a-Star Integrator (NESSI; Pietrow & Pastor Yabar, 2024) code to synthesize Sun-as-a-star emission from 20 solar flare observations using data with a limited FOV, obtained by the Swedish 1-m Solar Telescope (SST; Scharmer et al., 2003). This work will also serve as a pilot study on how the Paranal solar Espresso Telescope (POET; Leite et al., 2024; Santos et al., 2025) can observe flares within its limited FOV, ranging from 1″ to 60″, and therefore also explores what can be inferred from partially captured flares.
@@ 2 Data and methods
The data used in this work are a collection of observations spanning the period from 2011 to 2024, collected by various institutes for different scientific purposes. As a result, the observations use different line samplings, cadences, and pointing strategies. All sets are reduced by some version of the standard CRISPRED and SSTRED pipelines (de la Cruz Rodríguez et al., 2015; Löfdahl et al., 2021), but there are differences in the output products. Older cubes were saved in the legacy ”La Palma cube” format and contain polarimetric fringes and poor intensity calibration. These cubes were also not always rotated in such a way that solar north is up. In these cases, the rotation was performed manually based on the SST turret logs, intensity calibrations were performed, and the cubes were saved into the standard FITS format.
@@
[…]

### v2

@@ Abstract
Aims. We produced and analyzed synthetic full-disk spectra derived from 19 small area field-of-view optical observations of solar flares acquired by the Swedish 1-m Solar Telescope (SST) between 2011 and 2024. These were used to investigate what can and cannot be inferred about physical processes on the Sun from Sun-as-a-star observations.
Methods. The recently released Numerical Empirical Sun-as-a-Star Integrator (NESSI) code provides synthetic full-disk integrated spectral line emission based on smaller field-of-view input while accounting for center-to-limb variations and differential rotation. We used this code to generate pseudo-Sun-as-a-star spectra from the SST observations.
Results. We show that limited-area solar observations can be extrapolated to represent the full disk accurately in a manner close to what is achievable with Sun-as-a-star telescopes. Additionally, we identify nine spectral features, four of which are caused by instrumental effects. Most notably, we find a relation between the heliocentric angle of flares and the width of the excess emission left by them as well as a source of false positive coronal mass ejections-like signatures, and we defined an energy scaling law based on chromospheric line intensities that shows that the peak flare contrast roughly scales with the square root of the bolometric energy.
Conclusions. The presented method of creating pseudo-Sun-as-a-star observations from limited field-of-view solar observations allows for the accurate comparison of solar flare spectra with their stellar counterparts while allowing for the detection of signals at otherwise unachievable noise levels.
@@ 1 Introduction
Solar flares and the corresponding coronal mass ejections (CMEs) are the primary drivers of the most severe disruptions to the heliosphere (Cliver et al., 2022). When aimed at the Earth, they can cause widespread damage to electronics, blackouts, and disruptions in communications (e.g., Cole, 2003; Miteva et al., 2023). The energies of these flares are typically between and erg (Fletcher et al., 2011), which is still two orders of magnitude lower than the most powerful stellar flares. These extreme events are believed to reach energies as high as or erg, depending on whether one adopts the assumption of blackbody radiation (Maehara et al., 2012) or hydrogen recombination continuum (Simões et al., 2024) for energy fitting. While solar flares leave only faint (Woods et al., 2006; Kretzschmar, 2011) or at times no detectable (e.g., Pietrow et al., 2024a) signatures in the total solar irradiance, they can dramatically outshine their host stars when occurring on intrinsically dimmer M dwarfs (e.g., Zhan et al., 2019). They are also visible as dimmings in the corona and transition region (e.g., Hou et al., 2025; Veronig et al., 2025).
@@
Although such extreme events have been observed on other stars for years (e.g., Pietras et al., 2022; Namizaki et al., 2023; Bruno et al., 2024), stellar flares have only recently attracted significant attention in the context of exoplanet research. In part because they are sources of noise and transit time variation (e.g., Davenport et al., 2019; Howard et al., 2023) but also since they are agents of change in planetary environments, where it is believed that flares and their ejecta can permanently alter planetary atmospheres and thus their habitability (e.g., Pulkkinen, 2007; Airapetian et al., 2020; Temmer, 2021; Ilin et al., 2024).
@@
For this reason, it is important to understand stellar flares and what differentiate them from their solar counterparts. However, this is not straightforward, as solar flares are typically studied in high spectral and spatial resolution, while the opposite is true for stellar flares, which are primarily observed in white light, and only more recently with high-resolution spectrographs (e.g., Namekata et al., 2021; Mart´ınez González et al., 2022; Namizaki et al., 2023). As solar flares imprint only faint (Woods et al., 2006; Kretzschmar, 2011) or non-detectable (Pietrow et al., 2024a) white-light fluctuations on a disk-integrated scale, direct comparisons have so far only been possible within numerical models (e.g., Allred et al., 2005; Druett et al., 2024). However, this is no longer the case when looking at spectral lines where even relatively weak flares for stellar standards can leave an imprint on a disk-integrated spectrum (see Fig. 1).
@@
Namekata et al. (2021, 2022), Otsu et al. (2022), and Xu et al. (2022) pioneered a pseudo-Sun-as-a-star analysis technique for flares and CMEs by studying line-profile evolutions of spatially integrated spectra of a small region around the event. Here, a difference was seen between the imprint of the flare on the spectra and its position on the disk. Such a relation has already been known to exist between UV and X-ray radiation for solar flares (Woods et al., 2006) and has recently been suggested to also exist for spectral lines in the visible by Yu et al. (2025) through 1D flare simulations using the RADYN code (Allred et al., 2005). If confirmed in observations, such an effect could prove to be a powerful technique for determining the disk position of these events.
@@
Follow-up studies with such Sun-as-a-star telescopes are limited by their signal-to-noise ratio (S/N) and are primarily sensitive to very strong flares. On the other hand, there is a large collection of flare observations taken with high spatial and spectral resolution in the solar community that have excellent S/N but a limited FOV. Such observations can be processed in a way similar to that of other pseudo-Sun-as-a-star studies to obtain more statistics on flare behavior in a disk-integrated setting. However, we propose adding an additional step to more closely approximate true Sun-as-a-star flare observations. In this work, we apply the Numerical Empirical Sun-as-a-Star Integrator (NESSI; Pietrow & Pastor Yabar, 2024) code to synthesize Sun-as-a-star emission from 20 solar flare observations using data with a limited FOV, obtained by the Swedish 1-m Solar Telescope (SST; Scharmer et al., 2003). This work also serves as a pilot study on how the Paranal solar Espresso Telescope (POET; Leite et al., 2024; Santos et al., 2025) can observe flares within its limited FOV, ranging from 1″ to 60″, and therefore also explores what can be inferred from partially captured flares.
@@ 2 Data and methods
The data used in this work are a collection of observations spanning the period from 2011 to 2024, collected by various institutes for different scientific purposes. As a result, the observations use different line samplings, cadences, and pointing strategies. All sets are reduced by some version of the standard CRISPRED and SSTRED pipelines (de la Cruz Rodr´ıguez et al., 2015; Löfdahl et al., 2021), but there are differences in the output products. Older cubes were saved in the legacy ”La Palma cube” format and contain polarimetric fringes and poor intensity calibration. These cubes were also not always rotated in such a way that solar north is up. In these cases, the rotation was performed manually based on the SST turret logs, intensity calibrations were performed, and the cubes were saved into the standard FITS format.
@@
2.1 NESSI
The NESSI code (Pietrow & Pastor Yabar, 2024) was created to convert limited FOV solar observations into simulated full-disk Sun-as-a-star observations. This is done using accurate CLV measurements for a given wavelength range to create a solar disk on a polar grid. The line profiles on this grid are then shifted according to a differential rotation law, such as the one presented in Balthasar et al. (1986) and assumed to be constant for all lines. This grid is then integrated to create a quiet-Sun profile of the Sun-as-a-star. Flare observations can be injected by dividing their average spectra by a subfield of the NESSI quiet-Sun disk, matching the size and position of the SST FOV, after which the resulting spectrum is multiplied with the quiet-Sun disk. This makes it as if the SST flare had occurred on the full disk. The process is illustrated in Fig. 1. NESSI works in a similar way to the Spot Oscillation And Planet code (SOAP; Boisse et al., 2012; Dumusque et al., 2014) but relies fully on empirical data.
This approach of creating pseudo-Sun-as-a-star observations is similar to the one presented by Otsu et al. (2022), where small areas on the disk are integrated and normalized by quiet-Sun areas of similar size elsewhere on the disk. However, our approach allows for the preservation of intensities, thus showing how bright and detectable a flare would be in a true Sun-as-a-star observation such as the one presented in Pietrow et al. (2024a). This method is a digital analog to the observational strategy of the POET telescope, and many of the issues discussed in this work will be relevant to future observations with this telescope. The codes used in the follow-up analysis, the example images, and animations are publicly available at this repository111https://github.com/demichieldewilde/solar-flares, as well as included with this paper.
2.2 Data normalization
We determine the spectral contribution of each FOV by comparing the dataset to a quiet reference region contained within. For each observation, an area containing the ”quietest Sun” is manually selected inside one of the frames of the time series, away from the peak of the flare. Intensity data over this selected area are averaged to obtain the reference quiet-Sun spectral intensity222In the following, quantities representing spectral intensity are distinguished from specific intensities by the subscripts QR, FOV, and FD, which indicate area integrals over a Quiet-Sun Region, the Field of View, or the Full Disk, respectively. as a function of wavelength . The NESSI code is then used to generate a theoretical quiet-Sun profile at this disk location , and the two are scaled to calibrate the observational data to the NESSI disk data. The continuum wavelength is used where available, and otherwise the furthest wing wavelength points are used, as these are less sensitive to flares and other activity signals. The scaling constant is introduced such that the merit function
2.3 Contrast profiles
Therefrom the contrast profile can be defined as
2.4 Noise simulations
We made a case study based on observations of the X9.3 flare, where we compare our synthetic Sun-as-a-star contrast profiles with those of the same flare, which was also observed with the HARPS-N telescope (Pietrow et al., 2024a). Firstly, we match the 5-minute cadence by taking every 20th observation. We do not change the spectral resolution, as the resolution of SST/CRISP is very similar to the resolution of HARPS-N. The S/N of the observations are matched next, where we add noise based on the standard deviation of the HARPS-N continuum close to the Ca ii K and H lines. By matching the noise level of HARPS-N both photon and instrumental noise is accounted for. Finally, the resulting profile is multiplied by 2.5 to account for the fact that we capture only about 40% of this flare in our FOV.
The resulting contrast profiles for H and Ca ii K are shown in Fig. 3 in the top and bottom row, respectively. For both lines, the first column shows contrast profiles of the undegraded SST contrast profiles. The second column shows the same profiles degraded to the HARPS-N cadence and S/N, and the final column shows the corresponding HARPS-N profiles.
While the SST data was observed for a much shorter time, we find a strong similarity with the contrast profile shapes and intensities between minutes 10 and 20. After this time, the flare seemingly gets dimmer in the SST frame while it stays bright in the other data. This is consistent with the flaring activity inside and outside of the limited FOV of the SST observations, which captures only around 40% of the flare, i.e., toward the middle of flare’s lifetime, most of the bright ribbon feature lies outside of the SST FOV.
2.5 Differenced equivalent width
In our case, it takes the form
2.6 Residual analysis via Voigt profile subtraction
At each time frame , the contrast profile is modeled by the function
3 Results and discussion
3.1 Feature summary
3.1.1 Physical features
Our investigation finds the following features to originate from processes on the Sun, which can be interpreted physically. We distinguish them from other features.
Feature 1: Core intensity modulation
Feature 2: Redshifted emission
Several events display a transient, red‑shifted emission component that becomes most prominent at the flare peak and fades as the flare progresses. This signature is evident in Event 3 (Fig. 6; Ca ii 8542 Å, Ca ii K), Event 5 (Fig. 8; Ca ii 8542 Å), Event 6 (Fig. 9; Ca ii 8542 Å), Event 7 (Fig. 10; H, H), and Event 17 (Fig. 19; Ca ii 8542 Å, Ca ii K). We attribute this temporary red shift to down‑flowing, cool, condensed plasma that drains along flare loops into the flare ribbons, which are largely unobscured by overlying filaments (see the corresponding animations). This phenomenon is commonly referred to as coronal rain (see e.g., Pietrow et al., 2024b). All of these events lie at heliocentric directional cosines of (Table 2), a range that maximizes the line‑of‑sight projection of the flow velocity and thereby enhances the red‑shifted components. This interpretation is in line with the findings of Yu et al. (2025). A redshifted emission – referred to as red asymmetry – is evident in the H profile associated with a superflare on YZ Canis Minoris, reported by Namizaki et al. (2023, panel (d) in Fig. 3).
Feature 3: Imprints of filaments and flare loops
Feature 4: Temporal offset between and GOES peaks
3.1.2 Spurious features
We attribute these features primarily to instrumental and seeing-induced effects, which should be carefully considered in future studies involving a limited FOV. We highlight these features as they can be misinterpreted as physical features.
Feature 5: False quasi-periodic oscillations
Feature 6: False profile modulation and abrupt jumps
3.1.3 Detectability of CMEs
3.2 Individual event summaries
Event 1: An X9.3 flare on 2017 September 6
An X9.3 flare occurred in the active region NOAA 12673 on 2017-09-06 at 11:53 UT. Approximately 40% of the flare ribbon area was captured in the FOV of the SST, and therefore, we scaled the recorded flare emission by a factor of 2.5 before passing it onto NESSI as described in Section 2.3. The average value for the flare is 0.82. The flare was associated with a Halo CME with a recorded velocity of 1571 km s-1 and CME-driven shock that intercepted the Earth causing a significant space weather event (e.g., Yu et al., 2024). This flare was first described in (Quinn et al., 2019).
Event 2: An X8.2 flare on 2017 September 10
Event 3: An X2.2 flare on 2017 September 6
Event 4: An X1.5 flare on 2014 June 10
Event 5: An X1.0 flare on 2021 October 28
Event 6: An M3.0 flare on 2022 May 20
Event 7: An M2.5 flare on 2023 June 9
There is a blue-shifted absorption feature in the Ca ii 8542 Å and H line cores at times after minutes. This is a result of taking the reference frames for the contrast profile around 55-60 minutes during the gradual phase of the flare. As the ribbon activity decreases slowly, the ”contrast” therefore appears negative in the core at later times.
Event 8: An M1.9 flare on 2015 September 27
In this instance throughout the observations, there is an arcade of loops that overlaps the flare ribbon emission in the line of sight (Fig. 11, bottom right panel). From the line profiles and the cocoplot (particularly in Ca ii 8542 Å), one can infer there is a transfer of relatively cool material along these arched fibrils occurring predominantly from one ribbon at () = (785″, ″) toward the other ribbon at () = (780″, ″) producing a net emission showing a blue Doppler shift in the line core.
As a result, the contrast profiles are strongly impacted throughout the duration of this action. The line core minimum position shifts left, to the shorter or bluer wavelengths. Therefore, relative to the reference spectrum, the emission in the wing wavelength left of the center appears lower, and that to the right appears greater. Additionally, the peak in the line s is much later than the GOES peak in both cases. We believe that this is due to the cool material covering the flare ribbons.
When looking at a contrast profile or residual, a persistent net Doppler shift to one side generates a pair of enhanced and decreased emission on opposite side of the line core. Which side is enhanced and which decreased depends on whether the relevant material is behaving as an emission line (enhancement matches Doppler shift) or an absorption line (enhancement in opposite direction to Doppler shift). However, this signal is not associated with a related CME, despite exhibiting the CME spectral signature that stellar observers often look for (e.g., Namekata et al., 2021).
Event 9: An M1.8 flare on 2024 September 11
Event 10: An M1.1 flare on 2014 September 6
This flare shows some striking differences to the M1.9 flare (Event 8). Both have similar strengths in GOES and occur near the limb. However, the emission from this flare is exceptionally symmetric in both the contrast profiles and the Voigt profile residuals (see Fig. 13). In this instance, no overlying absorbing material is present to provide absorption with a significant net Doppler shift, which results in the highly symmetric profiles. This emphasizes the power of coronal rain, filaments, and fibrils in determining the net Doppler velocities of Sun-as-a-star flare observations of chromospheric lines, especially when contrasted with the flare ribbons themselves. The same misinterpretation can also occur in spatially resolved studies, where individual wavelength images are inspected without appropriate context, showing what emission originates from flare ribbon and what spectral features are linked to overlying material. Context can easily be provided by a COCOPLOT.
Exceptionally, the observation includes a rather long segment of preflare data, which includes a detected microflare at minutes. We also observe a 50-minute line center darkening preceding the flare which then turns into faint red- and blue-shifted emission. The other investigated flares do not include long enough preflare data to check whether this is a common emission pattern, more pre-flare time series are needed to answer this question.
Event 11: A C8.4 flare on 2016 May 14
The impulsive phase of the flare started 5 to 10 minutes before the start of the SST observation (Fig. 14). A small filament/prominence was present over the classic sigmoid 333sigmoid flare ribbons are S-shaped, not shaped. flaring area. The cool prominence material produces strong absorption in chromospheric lines where it covers the flaring area in the line of sight at times near the start of the SST observation. This is especially evident in H observations where the prominence material is highly opaque (Fig. 14, bottom right panel, strong blue and red features). In the first ten minutes of the SST observation (flare time minutes) the bring flare ribbon emission decreases, but there is also an evolution of the filament with up-and-downflows as well as the launch of some fraction as a CME. This reorganization and draining of the filaments clears the view of the flaring area leading to a clear second, more gradual ”peak” in the flare emission (Fig. 14, top right panel) which is particularly strong in H compared with Ca ii 8542 Å, again due to the high opacity of the filament in H.
The contrast profiles and Voigt residuals clearly show the effect of gravity acting on the overlying absorbing material in this filament from minutes from the deceleration of upflowing material (blue wing absorption of the spectral lines in Fig. 14, left and central panels), and the acceleration of downflowing material (subsequent red wing absorption), which manifests in an approximately linear drift of the Doppler velocities to longer wavelength over time. Back-of-the-envelope calculations of the linear drift in Doppler velocity ( to 60 km s-1 over a time of 25 minutes, viewed at angle give a constant acceleration of 0.210 km s-2. Given the errors in this calculation and the angle of the magnetic field with respect to the vertical, this action is consistent with draining under solar gravity (0.274 km s-2). There is a blue-shifted absorption feature seen in the Ca ii 8542 Å line core at times after minutes. As in Event 7, this is a result of taking the reference frames for the contrast profile during the gradual phase of the flare.
Event 12: A C5.6 flare on 2015 June 24
Event 13: A C4.1 flare on 2011 August 6
Event 14: A C3.6 flare on 2015 May 5
Event 15: A double C1.9 and C2.4 flare on 2022 June 26
The first flare happens partially out of the field of view and produces associated cool coronal loop material that does not become a CME. Instead the material falls back down, obscuring some part of the flare ribbons and creating a dual blue- and red-shifted absorption signatures. The second stronger flare starts toward the end of this process. From the pattern of the EWs alone it is not possible to tell whether this is one larger flare with emission that gets strongly absorbed near its start, or two smaller flares happening at relatively close times. In a true disk-integrated setting such an event could be mistaken for a single stronger flare if no X-ray or white light information is available. In Ca ii 8542 Åwe see dimming in the contrast profile core wavelengths after the flare in a fashion similar to Event 7.
Event 16: A C2.2 flare on 2023 June 18
Event 17: A C2.0 flare on 2019 May 6
Event 18: A C1.5 flare on 2013 June 30
The oscillations in intensity are once again seeing induced.
Event 19: A C1.2 flare on 2021 May 26
This event represents a particularly weak flare inside of a decaying active region that produced several more C-class flares in the following days before it rotated out of sight behind the limb. In this case, one primarily sees mildly tangled field lines linking the fragmented spots and pores of the decaying region. The flare occurs at the light bridge in the upper corner of the observation and starts off very delicately with several small micro-flare brightenings before a larger reconnection occurs around 7 minutes. Around 15 minutes the reconnection calms down before several new microflares occur around 20 minutes and toward the end of the time series.
3.3 Flare energy scaling law
| Line | |||
| Ca ii K | 0.995 | 0.430 | |
| Ca ii 8542 Å | 0.924 | 0.443 | |
| H | 0.847 | 0.417 | |
| Overall Fit | 0.744 | 0.547 |
4 Conclusions
In this work, we have used a unique collection of 20 solar flares observed with the SST (see Fig. 2) to obtain statistically relevant results. These limited-FOV time series were averaged and turned into approximations of disk-integrated spectra using the NESSI code (Pietrow & Pastor Yabar, 2024) which were shown to closely reproduce true Sun-as-a-star observations with certain caveats such as the incomplete capture of a flare, seeing variations, as well as FOV shifts, jumps and drifts. As noted, limited-FOV observations are prone to errors introduced by quiet-Sun areas adjoining sunspots or active regions. Our method mitigates these issues while exploiting the main advantage of a restricted FOV–namely, isolating the spectral influence of the flaring region itself. Partially covered flares therefore remain informative, because the behavior they reveal still contributes to the flare’s overall spectral evolution; however, caution is warranted, as unobserved portions of the event may introduce unaccounted-for effects.
Additionally, this work allowed us to validate these previous studies, as our flares are observed at a much higher spatial and temporal resolution. Perhaps unsurprisingly, we do not report a strong gain from the resolution increase, but we could temporally resolve much more of the flare structure. Additionally, the multi-line observations used in this work provided additional information that is often missing from single-line observations.
We believe that the spectral features described in this paper can help interpret unresolved stellar flares, as well as show that strong blue-shifted absorption does not necessarily indicate a CME. However, we also urge for more studies of this type with larger amounts samples of flares to help solidity these findings.
4.1 Data availability
Acknowledgements.
- Cole (2003) Cole, D. G. 2003, Space Sci. Rev., 107, 295
- de la Cruz Rodr´ıguez et al. (2015) de la Cruz Rodríguez, J., Löfdahl, M. G., Sütterlin, P., Hillberg, T., & Rouppe van der Voort, L. 2015, A&A, 573, A40
- Guo et al. (2023) Guo, J. H., Ni, Y. W., Zhong, Z., et al. 2023, ApJS, 266, 3
- Hou et al. (2025) Hou, Z., Tian, H., Yan, J., et al. 2025, A&A, 695, A12
- Kretzschmar (2011) Kretzschmar, M. 2011, A&A, 530, A84
- Libbrecht (2019) Libbrecht, T. 2019, PhD thesis, Stockholm University
- Liu et al. (2018) Liu, L., Cheng, X., Wang, Y., et al. 2018, ApJ, 867, L5
- Ma et al. (2024) Ma, Y. L., Lao, Q. H., Cheng, X., et al. 2024, ApJ, 966, 45
- Mart´ınez González et al. (2022) Martínez González, M. J., Asensio Ramos, A., González-Fernández, C., et al. 2022, MNRAS, 517, 744
- Otsu & Asai (2024) Otsu, T. & Asai, A. 2024, ApJ, 964, 75
- Pietrow (2022) Pietrow, A. G. M. 2022, PhD thesis, Stockholm University
- Temmer (2021) Temmer, M. 2021, Living Rev. Sol. Phys., 18, 4
- Warmuth & Mann (2016) Warmuth, A. & Mann, G. 2016, A&A, 588, A116
- Xu et al. (2022) Xu, Y., Tian, H., Hou, Z., et al. 2022, ApJ, 931, 76
- Yu et al. (2025) Yu, H. C., Hong, J., & Ding, M. D. 2025, A&A, 694, A315
- Yu et al. (2024) Yu, X. X., Zhang, S. N., Lu, H., et al. 2024, ApJ, 960, 85
Appendix A Event table
| No. | DateaaaaYYYY-MM-DD. | NOAAbbbbFrom SolarSoft Latest Events Archive. | ClassbbbbFrom SolarSoft Latest Events Archive. | ccccLinear Fit Speed from the SoHO / LASCO CME Catalog (Gopalswamy et al. 2009). | Flare Start / Peak / StopbbbbFrom SolarSoft Latest Events Archive. | SST Start / Stop | OverlapddddSpatial overlap primarily estimated between the CRISP FOV and the total flare area in AIA 1700. | eeeeThe coordinates represent the average estimated center of the flare within the CRISP and/or CHROMIS FOV, determined by excluding highly variable regions and focusing on consistently covered areas. | Lines | Cad. | |||